OpenStudy (anonymous):

PLZ HELP MEDAL .. Consider the sequence 16, –8, 4, –2, 1, ... a. Describe the pattern formed in the sequence. b. Find the next three terms.

OpenStudy (anonymous):

divide by -2 each time

OpenStudy (anonymous):

-.5, .25, -.125

OpenStudy (anonymous):

so -1/2, ...?

OpenStudy (anonymous):

what did u divide by -2 first?

jimthompson5910 (jim_thompson5910):

alternatively, you multiply by -1/2 each time so the next 3 terms are: -1/2, 1/4, -1/8 since 1*(-1/2) = -1/2 -1/2*(-1/2) =1/4 1/4*(-1/2) = -1/8

OpenStudy (anonymous):

divide by -2 each time -.5, .25, -.125 For example, divide the number 16 by -2, you get -8. Now divide -8 by -2 and you get 4. The pattern is hard to grasp bc negative number and positive number become flipped when multiplying or dividing

OpenStudy (anonymous):

ok got it thank you